# Integrating Neural Operators with Diffusion Models Improves Spectral Representation in Turbulence Modeling

### v1

@@ 4.2 Score-based diffusion models
$\hat{s}(k_{i},U_{NO},\sigma_{j})\simeq\nabla_{\hat{u}(k_{i})}\log p(\hat{u}(k_{i})|\hat{U}_{NO}(k_{i}),\sigma_{j})=-\frac{\hat{u}(k_{i})-\hat{U}_{NO}(k_{i})}{\sigma^{2}(k_{i},\sigma_{j})},$ (22)
where $\hat{u}(k_{i})$ is the Fourier transform of the current state at wave number $k_{i}$ and $\sigma^{2}(k_{i},\sigma_{j})$ is the variance of the distribution conditioned on the noise level $\sigma_{j}$ at wave number $k_{i}$. Although the specific relation between $\sigma^{2}(k_{i},\sigma_{j})$ and $k_{i}$ is problem specific, typically [55],
$\sigma^{2}(k_{i},\sigma_{j})\propto\frac{1}{k_{i}^{p}}+\sigma_{j}^{2},$ (23)
where $p$ is a positive constant based on the data. Hence, as $k_{i}\rightarrow 0$, $\sigma^{2}(k_{i},\sigma_{j})$ increases and $\frac{\hat{u}(k_{i})-\hat{U}_{NO}(k_{i})}{\sigma^{2}(k_{i},\sigma_{j})}\rightarrow 0$. This implies a weak gradient value signifying smaller steps. Now if we look at the frequency domain of the Langevin dynamics, from Eqn 15
$\hat{u}_{j+1}=\hat{u}_{j}(k_{i})-\frac{\varepsilon}{2}\frac{\hat{u}(k_{i})-\hat{U}_{NO}(k_{i})}{\sigma^{2}(k_{i},\sigma_{j})}+\sqrt{\varepsilon}\hat{z}_{j}(k_{i}),$ (24)
it can be seen that for smaller $k_{i}$ the above equation changes to
$\hat{u}_{j+1}\simeq\hat{u_{j}}(k_{i})+\sqrt{\varepsilon}\hat{z}_{j}(k_{i}).$ (25)
This signifies that the Langevin dynamics updates are dominated by the noise term $\sqrt{\varepsilon}\hat{z}_{j}(k_{i})$ instead of the gradient driven term and hence leads to ineffective refinement of the low wave numbers. Hence, in an intermediate step during the diffusion process when the updates are dominated by the noise term, the energy spectrum can be given as
$E_{U_{D,j+1}}(k_{i})=\frac{|\hat{u}_{j}(k_{i})|^{2}}{N}+\frac{\varepsilon|\hat{z}_{j}(k_{i})|^{2}}{N}+\frac{2\sqrt{\varepsilon}\Xi[\hat{u}_{j}(k_{i})\hat{z}^{\ast}_{j}(k_{i})]}{N},$ (26)
where $\Xi[.]$ represents the real part of the complex number and $\hat{z}^{\ast}$ the complex conjugate of $\hat{z}$. Due to weaker gradients, the term $|\hat{u}_{j}(k_{i})|^{2}$ grows slowly and $\varepsilon|\hat{z}_{j}(k_{i})|^{2}$ may dominate, which might lead to flatter energy spectrum at low wave number and may not accurately represent the large scale features. Therefore, as as $k_{\xi}$ increases $\left|u(\mathbf{X},t)-(U_{NO}(\mathbf{X},t)+U_{D}(\mathbf{X},t))\right|\rightarrow 0$, or in other words for the proposed method to work effectively, the $U_{NO}(\mathbf{X},t)$ should be at-least be able to capture the mean flow, in order for the diffusion model to be effectively helping in adding the high frequency features to the predicted mean flow.
@@ Acknowledgements
Funding: This work was supported by the MURI grant FA9550-20-1-0358 and the the ONR Vannevar Bush Faculty Fellowship (N0001422-1-2795).
Author Contributions:
V.O.: Conceptualization, Methodology, Software, Formal Analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing and Visualization.
A.B.: Conceptualization, Methodology, Software, Formal Analysis, Investigation, Writing - Original Draft and Writing - Review & Editing
Z.Z.: Software, Validation, Formal Analysis, Investigation and Visualization
G.E.K.: Validation, Formal Analysis, Resources, Writing - Review & Editing, Supervision, Project administration and Funding acquisition
@@
Data and Materials Availability: The data for Kolmogorov flow was simulated using the publicly available pseudo-spectral solver provided in FNO’s repository. The buoyancy-driven transport data was from PDEArena. The LES datasets for turbulent airfoil and jet were downloaded from link. The Schlieren velocimetry dataset can be found here. The code developed in this study will be made available after publication in our GitHub repository.
@@ Appendix A POD Analysis
In the main manuscript, we investigate the energy spectra with respect to the wavenumber corresponding to the neural operator and diffusion model conditioned on the neural operator, computed on each time step separately. Here we perform a detailed proper orthogonal decomposition (POD) analysis by performing singular value decomposition on 10 consecutive timesteps of the trajectory, and plot the eigen energy spectrum. Additionally, we also visualize the corresponding eigen modes/basis required for accurate reconstruction of the aforementioned timesteps. The POD analysis for Kolmogorov flow Figure A.1, buoyancy driven transport Figure A.2, turbulent airfoil Figure A.3, 3D turbulent jet Figure A.4 and schlieren velocimetry Figure 8 are shown below.
Fig.  A.1: POD analysis of Kolmogorov flow. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator (FNO) and diffusion model conditioned on the neural operator with the ground truth solution. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.
Fig.  A.2: POD analysis of buoyancy-driven flow. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator (UNet) and diffusion model conditioned on the neural operator with the ground truth solution. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.
Fig.  A.3: POD analysis of a turbulent airfoil. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator (MATCHO) and diffusion model conditioned on the neural operator with the ground truth solution. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.
Fig.  A.4: POD analysis of a 3D jet We perform POD on the angular velocity component at $\theta=0$. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator (MATCHO) and diffusion model conditioned on the neural operator with the ground truth solution. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.

### v2

@@ 4.2 Score-based diffusion models
$\hat{s}(\mathbf{X},\mathbf{\mathcal{G}},\sigma_{j})\simeq\nabla_{\hat{X}}\log p(\hat{\mathbf{X}}(k)|\hat{\mathcal{G}}(k))=-\frac{\hat{\mathbf{X}}(k)-\hat{\mathcal{G}}(k)}{var(k,\sigma_{j})},$ (22)
where $\hat{\mathbf{X}}(k)$ is the Fourier transform of the current state at wave number $k$ and $var(k,\sigma_{j})$ is the variance of the distribution conditioned on the noise level $\sigma_{j}$ at wave number $k$. Although the specific relation between $var(k,\sigma_{j})$ and $k$ is problem specific, typically [69],
$var(k,\sigma_{j})\propto\frac{1}{k^{p}}+\sigma_{j}^{2},$ (23)
where $p$ is a positive constant based on the data. Hence, as $k\rightarrow 0$, $var(k,\sigma_{j})$ increases and $\frac{\hat{u}(k)-\hat{\mathcal{G}}(k)}{var(k,\sigma_{j})}\rightarrow 0$. This implies a weak gradient value signifying smaller steps. Now if we look at the frequency domain of the Langevin dynamics, from Eqn 15
$\hat{\mathbf{X}}_{j+1}=\hat{\mathbf{X}}_{j}(k)-\frac{\varepsilon}{2}\frac{\hat{\mathbf{X}}(k)-\hat{\mathcal{G}}(k)}{var(k,\sigma_{j})}+\sqrt{\varepsilon}\hat{z}_{j}(k),$ (24)
it can be seen that for smaller $k$ the above equation changes to
$\hat{\mathbf{X}}_{j+1}\simeq\hat{\mathbf{X}_{j}}(k)+\sqrt{\varepsilon}\hat{z}_{j}(k).$ (25)
This signifies that the Langevin dynamics updates are dominated by the noise term $\sqrt{\varepsilon}\hat{z}_{j}(k)$ instead of the gradient driven term and hence leads to ineffective refinement of the low wave numbers. Hence, in an intermediate step during the diffusion process when the updates are dominated by the noise term, the energy spectrum can be given as
$E_{s_{\theta_{D}},j+1}(k)=\frac{|\hat{\mathbf{X}}_{j}(k)|^{2}}{N}+\frac{\varepsilon|\hat{z}_{j}(k)|^{2}}{N}+\frac{2\sqrt{\varepsilon}\Xi[\hat{\mathbf{X}}_{j}(k)\hat{z}^{\ast}_{j}(k)]}{N},$ (26)
where $\Xi[.]$ represents the real part of the complex number and $\hat{z}^{\ast}$ the complex conjugate of $\hat{z}$. Due to weaker gradients, the term $|\hat{\mathbf{X}}_{j}(k)|^{2}$ grows slowly and $\varepsilon|\hat{z}_{j}(k)|^{2}$ may dominate, which might lead to flatter energy spectrum at low wave number and may not accurately represent the large scale features.
Therefore, as as $k_{c}$ increases $\left|u(\mathbf{x},t)-s_{\theta_{D}}(\mathbf{X},\mathbf{\mathcal{G}},\sigma))\right|\rightarrow 0$, or in other words for the proposed method to work effectively, the $\mathbf{\mathcal{G}}$ should be at-least be able to capture the mean flow, in order for the diffusion model to be effectively helping in adding the high frequency features to the predicted mean flow.
@@ Acknowledgements
V.O. acknowledges Alan John Varghese, and Dr. Andreas Euan Robertson for their insightful suggestions and valuable discussions at various stages of this project.
Funding: This work was supported by the MURI grant (FA9550-20-1-0358), the ONR Vannevar Bush Faculty Fellowship (N0001422-1-2795)and the U.S. Department of Energy, Advanced Scientific Computing Research program, under the Scalable, Efficient and Accelerated Causal Reasoning Operators, Graphs and Spikes for Earth and Embedded Systems (SEA-CROGS) project, (DE-SC0023191).
Author Contributions: V.O.: Conceptualization, Methodology, Software, Formal Analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing and Visualization. A.B.: Conceptualization, Methodology, Software, Formal Analysis, Investigation, Writing - Original Draft and Writing - Review & Editing Z.Z.: Software, Validation, Formal Analysis, Investigation, Visualization, Writing - Original Draft and Writing - Review & Editing G.E.K.: Validation, Formal Analysis, Resources, Writing - Review & Editing, Supervision, Project administration and Funding acquisition
@@
Data and Materials Availability: The data for Kolmogorov flow was simulated using the publicly available pseudo-spectral solver provided in FNO’s repository. The buoyancy-driven transport data was from PDEArena. The LES datasets for turbulent airfoil and jet were downloaded from link. The Schlieren velocimetry dataset can be found here. The code developed in this study will be made available after publication in our GitHub repository - https://github.com/vivekoommen/NeuralOperator_DiffusionModel.
@@ Appendix A POD Analysis
In the main manuscript, we investigate the energy spectra with respect to the wavenumber corresponding to the neural operator and diffusion model conditioned on the neural operator, computed on each time step separately. Here we perform a detailed proper orthogonal decomposition (POD) analysis by performing singular value decomposition on 10 consecutive timesteps of the trajectory, and plot the eigen energy spectrum. Additionally, we also visualize the corresponding eigen modes/basis required for accurate reconstruction of the aforementioned timesteps. The POD analysis for Kolmogorov flow Figure 10, buoyancy driven transport Figure A.1, turbulent airfoil Figure A.2, 3D turbulent jet Figure A.3 and schlieren velocimetry Figure A.4 are shown below.
Fig.  A.1: POD analysis of buoyancy-driven flow. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator (UNet) and diffusion model conditioned on the neural operator with the ground truth solution. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.
Fig.  A.2: POD analysis of a turbulent airfoil. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator (TC-UNet) and diffusion model conditioned on the neural operator with the ground truth solution. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.
Fig.  A.3: POD analysis of a 3D jet We perform POD on the angular velocity component at $\theta=0$. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator (TC-UNet) and diffusion model conditioned on the neural operator with the ground truth solution. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.
Fig.  A.4: POD analysis of Schlieren velocimetry. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator and diffusion model conditioned on neural operator (TC-UNet) with the true solution obtained from Schlieren velocimetry. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.